1. A uniform electric field E = 6500 N/C passing through a flat square area

A = 10m2


. What is the electric flux?


1
Expert's answer
2022-03-24T11:01:19-0400

The electric flux

"\phi=EA\\
=\rm 6500\: N/C\times 10\: m^2=65000\: N\cdot m^2/C"